OpenStudy (anonymous):

34/6 = x/6 X= what ? A) 15.75 B) 17.55 C) 40.80 D) 28.90

OpenStudy (anonymous):

i got 5.66667 when i divided 34 into 6. but nothing i multiply by the 2nd equation equals that number. im confused /: @dpaInc @jim_thompson5910

jimthompson5910 (jim_thompson5910):

cross multiply to go from 34/6 = x/6 to 34*6 = 6x

OpenStudy (anonymous):

wouldnt it be 34 ? i dont get this problem at all

jimthompson5910 (jim_thompson5910):

yes it would be 34, so idk what's wrong with it either